I can't remember how to do this type of problem, someone please help me I will be so thankful. Find three consecutive intergers whose sum is 93.
Brina

Answer by Annabelle1(69)

let your three consecutive numbers be x , x+1 and x+2
the sum of these are x+(x+1)+(x+2) which simplifies to 3x+3
This is equal to 93 so 3x+3=93
minus 3 from both sides
3x+3-3=93-3
3x=90
divivde both sides by 3
x=30
The first integer is 30 so the following two are 31 and 32